Page 1 of 4

2018 Operation

Posted: Sun Mar 11, 2018 12:33 am
by dukwon
The LHC has been handed over to OP, and powering tests should begin Monday.

If you haven't done so already, you can catch up with the slides from Chamonix while waiting for first beams: https://indico.cern.ch/event/705545/att ... h_2018.pdf

Here's a nice article from the Bulletin:
https://home.cern/cern-people/updates/2 ... ibernation

2018 schedule:
https://beams.web.cern.ch/sites/beams.w ... e_2018.pdf

To summarise:
  • 60 /fb for ATLAS and CMS
  • 2 /fb for LHCb
  • 131 days of physics
  • 4 weeks of heavy ions
  • 1 week of desqueezed beams
  • Levelling at IP1&5 to a pile-up of 55 to 60 if needed
  • Continuous crossing-angle anti-levelling to improve lumi lifetime
  • 25—27 cm β* with ATS towards the end of the fill
  • 20 days of MD
Here's hoping for a nice physics production year to round off Run 2.

Re: 2018 Operation

Posted: Tue Mar 20, 2018 10:04 am
by dukwon
The short-term schedule is currently being driven by the intervention on the CMS pixel detector.

Of course things are subject to change, but here are some planned milestones

22nd March: first dry run with handshakes and mode changes
27th March: TED shots
29th March: vacuum valves open around CMS, possibly test of beam and injection permits
1st April: first beams
23rd April: first stable beams

Re: 2018 Operation

Posted: Thu Mar 29, 2018 9:13 pm
by jmc2000
Late February LHCC meeting:
https://indico.cern.ch/event/692482/con ... Status.pdf

Looks as if the number of days for the p-p run has been reduced from 131 to 124 to accommodate special runs:
https://beams.web.cern.ch/sites/beams.w ... e_2018.pdf

jmc2000

Re: 2018 Operation

Posted: Fri Mar 30, 2018 10:45 am
by dukwon
Both beams successfully circulating

https://i.imgur.com/3ef5c2g.png

Re: 2018 Operation

Posted: Tue Apr 10, 2018 10:02 am
by dukwon
The recommissioning is going very smoothly. Collision-finding should already take place this week, and first stable beams with 3 bunches could be as early as Monday 16th.

There will be a scrubbing run either next week or the one after.

Re: 2018 Operation

Posted: Wed Apr 11, 2018 7:04 pm
by dukwon
First collisions (not stable beams) could be tomorrow mid-morning if loss maps and TCDI validation are performed during the night.

Re: 2018 Operation

Posted: Thu Apr 12, 2018 7:28 am
by DCWhitworth
Thanks for keeping us all updated. I think I shall be having Page 1 and the LHC Announcer open on my work PC today :-)

Re: 2018 Operation

Posted: Thu Apr 12, 2018 9:22 am
by dukwon
dashboard0.png
We have collisions!

Re: 2018 Operation

Posted: Sat Apr 14, 2018 11:21 am
by dukwon
ADT tests (necessary for train injections and abort gap cleaning) have been given priority over first stable beams w/ 3b, so the latter has been pushed back to Tuesday.

Re: 2018 Operation

Posted: Mon Apr 16, 2018 6:47 pm
by dukwon
Latest news is that we could see stable beams after tomorrow's morning meeting (so in about 12 hours or more)

Re: 2018 Operation

Posted: Tue Apr 17, 2018 11:02 am
by dukwon
Stable beams with 3 bunches per beam!
lhc1.png

Re: 2018 Operation

Posted: Tue Apr 17, 2018 11:10 am
by dukwon
Screenshot from 2018-04-17 13-07-37.png
3DTower.png
Screenshot from 2018-04-17 13-07-00.png

Re: 2018 Operation

Posted: Sat Apr 21, 2018 11:37 am
by dukwon
We've had 3 fills with 75b, where the third was with BCMS beams. I believe the next step will be 300b.
Currently scrubbing with 2605b. More scrubbing planned on Monday.

Re: 2018 Operation

Posted: Mon Apr 23, 2018 8:09 am
by dukwon
We've had 2 fills with 339b. Scrubbing all of today until tomorrow morning. Then a third 339b fill with 48 bpi. Might switch to 600b Wednesday afternoon.

Re: 2018 Operation

Posted: Tue Apr 24, 2018 7:29 pm
by dukwon
Good news: the intensity ramp is about 7–10 days ahead of schedule, although some of that extra time may be spent on extra MDs.

Bad news: the problem at 16L2 has not gone away, and seems to be identical to last year. It was responsible for 5 dumps during the scrubbing run yesterday.

As a reminder: 16L2 stands for cell 16, left of point 2. Throughout 2017 there were a number of UFO-like events located between a quadrupole and a dipole in 16L2, which at large beam intensities would initiate a steady increase in losses over the next few turns until losses at IP7 caused a dump. It is believed that these were caused by air that had accidentally entered the pipe during the pump-down after EYETS and frozen to the beam screen. Warming up the beam screen in August made the problem worse, and the eventual solution was to operate with "8b4e" beams, referring to gaps of 100ns (4 empty) after every 8 bunches in a train. This mitigated the losses but unfortunately reduced the number of bunches from 2556 to 1920.

During YETS, 16L2 was warmed to 90 K and approximately 8.5 g of gas was extracted, which, when analysed, was found to be consistent with air. However, as we saw yesterday, this doesn't seem to have eliminated the problem.

Here's the discussion in last year's thread: http://www.lhcportal.com/Forum/viewtopi ... =45#p12568

Here are some LHC reports from last year which cover the 16L2 issue and 8b4e operation:
https://home.cern/cern-people/updates/2 ... ng-nothing
https://home.cern/cern-people/updates/2 ... tion-holes

Here are some relevant (publicly-available) slides from https://indico.cern.ch/event/705545/ and https://indico.cern.ch/event/664048/con ... s/2718791/
Screenshot from 2018-04-24 21-13-24.png
Screenshot from 2018-04-24 21-06-19.png
Screenshot from 2018-04-24 21-06-53.png
Screenshot from 2018-04-24 21-08-10.png
Screenshot from 2018-04-24 21-08-32.png
Screenshot from 2018-04-24 20-50-13.png
Screenshot from 2018-04-24 20-50-15.png
Screenshot from 2018-04-24 20-50-17.png
Screenshot from 2018-04-24 20-50-19.png